# An Inexact Bregman Proximal Difference-of-Convex Algorithm with Two Types of Relative Stopping Criteria

### v1

@@ 5.1 The $\ell_{1-2}$ regularized least squares problem
and choose the quadratic kernel function $\phi(\bm{x}):=\frac{1}{2}\|\bm{x}\|^{2}$. It is easy to verify that, with such choices, $(f,\,\phi)$ is 0-smooth adaptable restricted on $\mathbb{R}^{n}$. Moreover, we assume that the $A$ in (5.1) does not have zero columns so that $F_{\text{reg}}$ is level-bounded; see [21, Example 4.1(b)] and [50, Lemma 3.1]. Thus, our iBPDCA is applicable and the associated subproblem at the $k$-th iteration ($k\geq 0$) takes the following form:
@@
where $\bm{z}\in\mathbb{R}^{m}$ is the dual variable and $\bm{v}_{k}(\bm{z}):=\gamma_{k}^{-1}\bm{\xi}^{k}+\bm{x}^{k}-\gamma_{k}^{-1}A^{\top}\bm{z}$. From the property of the Moreau envelope of $\lambda\gamma_{k}^{-1}\|\cdot\|_{1}$ (see, e.g., [6, Proposition 12.29]), we see that $\Psi_{k}^{\text{reg}}$ is strongly convex and continuously differentiable with the gradient
@@
It is clear that all elements in ^$\widehat{\partial}^{2}\Psi_{k}^{\text{reg}}(\bm{z})$ are positive definite. We are now ready to present the Ssn method for solving equation (5.5) in Algorithm 2 and refer readers to [20, Theorem 3.6] for its convergence results.
@@
We next show that our inexact stopping criteria (SC1) and (SC2) can be achieved through some appropriate manipulations based the dual sequence generated by the Ssn method. Specifically, at the $k$-th iteration, we apply the Ssn method for solving equation (5.5), which generates a dual sequence $\{\bm{z}^{k,t}\}$. Let
@@ 5.1.2 Comparison results
We also include a well-known proximal difference-of-convex algorithm with extrapolation (pDCAe) [^1] The Matlab codes of the pDCAe for solving (5.1) are available at https://www.polyu.edu.hk/ama/profile/pong/pDCAe_final_codes/. [47] in comparisons. The pDCAe for solving (5.1) is given as follows: choose proper extrapolation parameters $\{\beta_{k}\}$, let $\bm{x}^{-1}=\bm{x}^{0}$, and then at the $k$-th iteration,
$&\mathrm{Take}~{}\mathrm{any}~{}\bm{\xi}^{k}\in\partial P_{2}(\bm{x}^{k})~{}\mathrm{and}~{}\mathrm{compute}\\ &\bm{y}^{k}=\bm{x}^{k}+\beta_{k}(\bm{x}^{k}-\bm{x}^{k-1}),\\ &\bm{x}^{k+1}=\mathop{\mathrm{argmin}}\limits_{\bm{x}}\left\{\langle A^{\top}(A\bm{y}^{k}-\bm{b})-\bm{\xi}^{k},\,\bm{x}\rangle+\frac{L_{A}}{2}\|\bm{x}-\bm{y}^{k}\|^{2}+\lambda\|\bm{x}\|_{1}\right\},\end{aligned}\right.$
where $L_{A}:=\lambda_{\max}(A^{\top}A)$ is the largest eigenvalue of $A^{\top}A$. We follow [47, Section 5] to choose the extrapolation parameters $\{\beta_{k}\}$ and perform both the fixed restart and the adaptive restart strategies.
@@
In the following experiments, we choose $\lambda\in\{0.01,0.1,1,10\}$ and consider $(m,n,s)=(100i,1000i,20i)$ for $i=2,5,10,15,20$. For each triple $(m,n,s)$, we follow [47, Section 5] to randomly generate a trial as follows. First, we generate a matrix $A\in\mathbb{R}^{m\times n}$ with i.i.d. standard Gaussian entries. We then choose a subset $\mathcal{S}\subset\{1,\cdots,n\}$ of size $s$ uniformly at random and generate an $s$-sparse vector $\bm{x}_{\text{orig}}\in\mathbb{R}^{n}$, which has i.i.d. standard Gaussian entries on $\mathcal{S}$ and zeros on the complement set $\mathcal{S}^{c}$. Finally, we generate the vector $\bm{b}\in\mathbb{R}^{m}$ by setting $\bm{b}=A\bm{x}_{\text{orig}}+0.01\cdot\widehat{\bm{n}}$, where $\widehat{\bm{n}}\in\mathbb{R}^{m}$ is a random vector with i.i.d. standard Gaussian entries.
@@
The average computational results for each triple $(m,n,s)$ from 20 instances are presented in Tables 1 and 2. From the results, one can observe that both iBPDCA-SC1 and iBPDCA-SC2 exhibit superior numerical performance compared to pDCAe, especially when the regularization parameter $\lambda$ is small or the problem size is large. For example, for the cases where $\lambda\leq 1$ and $n\geq 10000$, both iBPDCA-SC1 and iBPDCA-SC2 are consistently about 5 times faster than pDCAe, while achieving better objective function values. Moreover, one can also see that iBPDCA-SC1 and iBPDCA-SC2 exhibit the similar performance. This is indeed expected since they essentially employ the same framework, but differ in their stopping criteria for solving the subproblems. With the kernel function chosen as $\phi(\bm{x})=\frac{1}{2}\|\bm{x}\|^{2}$ in this part of experiments, the cost of computing the associated Bregman distance becomes negligible, rendering the effort required to verify inequalities (5.6) and (5.7) almost identical. Therefore, the CPU time of iBPDCA-SC1 and iBPDCA-SC2 can be similar when they undergo a similar number of total Ssn iterations.
@@
We also test all the methods with two instances $(A,\bm{b})$ obtained from the data sets mpg7 and gisette in the UCI data repository [17]. Here, the mpg7 data set is an expanded version of its original version, with the last digit signifying that an order 7 polynomial is used to generate the basis functions; see [20, Section 4.1] for more details. For mpg7, the matrix $A$ is of size $392\times 3432$ with $\lambda_{\max}(A^{\top}A)\approx 1.28\times 10^{4}$. For gisette, the matrix $A$ is of size $1000\times 4971$ with $\lambda_{\max}(A^{\top}A)\approx 3.36\times 10^{6}$. For each data set, we choose $\lambda=\lambda_{c}\|A^{\top}\bm{b}\|_{\infty}$ with $\lambda_{c}\in\{10^{-3},10^{-4}\}$. Figure 1 shows the numerical results of pDCAe, iBPDCA-SC1 and iBPDCA-SC2, where we plot the normalized objective value against the computational time. The normalized objective value is computed by $\frac{F_{\text{reg}}(\bm{x}^{k})-F_{\text{reg}}^{\min}}{F_{\text{reg}}(\bm{x}^{0})-F_{\text{reg}}^{\min}}$, where $F_{\text{reg}}^{\min}$ denotes the minimum of the terminating objective values obtained among all methods. From the results, we see that pDCAe exhibits a slower rate of reduction in the objective value, potentially attributed to the large Lipschitz constant associated with the real data. In contrast, our iBPDCA-SC1 and iBPDCA-SC2 continue to demonstrate superior performance on these two real data sets.
@@
$\quad&\hskip 79.6678pt{\text{pDCAe}}\\ &\hskip 91.04872pt\Downarrow\\ &\min\limits_{\bm{x}\in\mathbb{R}^{n}}~{}{\overbrace{{\lambda\|\bm{x}\|_{1}}}^{P_{1}(\bm{x})}\,-{\overbrace{{\lambda\|\bm{x}\|}}^{P_{2}(\bm{x})}}+\,{\overbrace{{\frac{1}{2}\|A\bm{x}-\bm{b}\|_{2}^{2}}}^{f(\bm{x})}}}\end{aligned}\quad~{}~{}vs~{}~{}\quad\begin{aligned} &\hskip 79.6678pt{\text{iBCDA}}\\ &\hskip 91.04872pt\Downarrow\\ &\min\limits_{\bm{x}\in\mathbb{R}^{n}}~{}{\overbrace{{\lambda\|\bm{x}\|_{1}+\frac{1}{2}\|A\bm{x}-\bm{b}\|_{2}^{2}}}^{P_{1}(\bm{x})}\,-\overbrace{{\lambda\|\bm{x}\|}}^{P_{2}(\bm{x})}\,+\,\overbrace{{0}}^{f(\bm{x})}}.\end{aligned}$
These two decomposition schemes result in different algorithms with different numerical performances. Since pDCAe also approximates the smooth convex part $\frac{1}{2}\|A\bm{x}-\bm{b}\|^{2}$ by a quadratic majorant, its performance would be significantly influenced by the Lipschitz constant $L_{A}:=\lambda_{\max}(A^{\top}A)$. This sensitivity to the Lipschitz constant is indeed a well-recognized limitation of first-order methods and may lead to the inferior performance of pDCAe for solving large-scale problems, as observed in Tables 1&\&&2 and Figure 1. This is because, for a randomly generated matrix $A$, the largest eigenvalue of $A^{\top}A$ tends to increase with the size, and for the data sets mpg7 and gisette, the largest eigenvalue of $A^{\top}A$ is around $1.28\times 10^{4}$ and $3.36\times 10^{6}$, respectively. In contrast, our iBPDCA can sidestep the majorization of the least squares term and choose instead to solve the subproblem via a highly efficient second-order Ssn method. The results in Tables 1&\&&2 and Figure 1 illustrate the promising performance of this algorithmic framework. But we should emphasize that these achievements are made possible by allowing the inexact minimization of the subproblem. This also highlights the motivation for developing an efficient inexact algorithmic framework in this work.
@@
(200,2000,40) pDCAe 2.73e-1 30000 2.12 0.02 2.54e+0 8353 0.59 0.03
iBPDCA-SC1 2.54e-1 120 (405) 0.63 0.02 2.54e+0 30 (170) 0.30 0.03
iBPDCA-SC2 2.54e-1 120 (414) 0.63 0.02 2.54e+0 30 (173) 0.30 0.03
(500,5000,100) pDCAe 8.12e-1 30000 10.83 0.17 6.92e+0 24732 8.88 0.16
iBPDCA-SC1 6.92e-1 161 (599) 4.75 0.17 6.92e+0 35 (259) 2.47 0.16
iBPDCA-SC2 6.92e-1 161 (612) 4.79 0.17 6.92e+0 35 (263) 2.47 0.16
(1000,10000,200) pDCAe 1.88e+0 30000 134.67 1.90 1.50e+1 30000 134.72 1.90
iBPDCA-SC1 1.48e+0 242 (898) 29.48 1.90 1.48e+1 50 (379) 15.46 1.90
iBPDCA-SC2 1.48e+0 242 (914) 29.65 1.90 1.48e+1 50 (382) 15.27 1.90
(1500,15000,300) pDCAe 2.99e+0 30000 359.23 5.04 2.31e+1 30000 359.53 5.03
iBPDCA-SC1 2.24e+0 299 (1072) 80.92 5.04 2.24e+1 59 (439) 41.27 5.03
iBPDCA-SC2 2.24e+0 299 (1088) 81.28 5.04 2.24e+1 59 (444) 41.06 5.03
(2000,20000,400) pDCAe 4.12e+0 30000 651.03 9.31 3.13e+1 30000 651.14 9.31
iBPDCA-SC1 3.00e+0 367 (1295) 187.79 9.31 3.00e+1 73 (518) 91.52 9.31
iBPDCA-SC2 3.00e+0 367 (1312) 187.91 9.31 3.00e+1 73 (524) 91.17 9.31
@@
(200,2000,40) pDCAe 2.53e+1 515 0.04 0.02 2.44e+2 153 0.01 0.02
iBPDCA-SC1 2.53e+1 9 (50) 0.10 0.02 2.44e+2 8 (32) 0.08 0.02
iBPDCA-SC2 2.53e+1 9 (50) 0.10 0.02 2.44e+2 8 (32) 0.08 0.02
(500,5000,100) pDCAe 6.90e+1 2633 1.00 0.16 6.80e+2 237 0.10 0.16
iBPDCA-SC1 6.90e+1 12 (116) 1.20 0.16 6.80e+2 7 (45) 0.54 0.16
iBPDCA-SC2 6.90e+1 12 (117) 1.21 0.16 6.80e+2 7 (45) 0.54 0.16
(1000,10000,200) pDCAe 1.48e+2 8547 38.41 1.90 1.47e+3 401 1.84 1.91
iBPDCA-SC1 1.48e+2 15 (185) 9.00 1.90 1.47e+3 7 (69) 3.79 1.91
iBPDCA-SC2 1.48e+2 15 (184) 8.90 1.90 1.47e+3 7 (68) 3.77 1.91
(1500,15000,300) pDCAe 2.24e+2 13670 163.91 5.04 2.23e+3 777 9.41 5.04
iBPDCA-SC1 2.24e+2 17 (225) 25.18 5.04 2.23e+3 8 (93) 12.13 5.04
iBPDCA-SC2 2.24e+2 17 (226) 25.06 5.04 2.23e+3 8 (93) 12.08 5.04
(2000,20000,400) pDCAe 2.99e+2 20631 448.03 9.33 2.98e+3 1685 36.88 9.30
iBPDCA-SC1 2.99e+2 20 (264) 55.59 9.33 2.98e+3 9 (129) 30.92 9.30
iBPDCA-SC2 2.99e+2 20 (265) 55.19 9.33 2.98e+3 9 (129) 30.73 9.30
@@
Figure 1: Numerical results of pDCAe, iBPDCA-SC1 and iBPDCA-SC2 for solving the $\ell_{1-2}$ regularized least squares problem on mpg7 and gisette from the UCI data repository.

### v2

@@ 5.1 The $\ell_{1-2}$ regularized least squares problem
and choose the quadratic kernel function $\phi(\bm{x}):=\frac{1}{2}\|\bm{x}\|^{2}$. It is easy to verify that, with such choices, $(f,\,\phi)$ is 0-smooth adaptable restricted on $\mathbb{R}^{n}$. Moreover, we assume that the $A$ in (5.1) does not have zero columns so that $F_{\text{reg}}$ is level-bounded; see [25, Example 4.1(b)] and [53, Lemma 3.1]. Thus, our iBPDCA is applicable and the associated subproblem at the $k$-th iteration ($k\geq 0$) takes the following form:
@@
where $\bm{z}\in\mathbb{R}^{m}$ is the dual variable and $\bm{v}_{k}(\bm{z}):=\gamma_{k}^{-1}\bm{\xi}^{k}+\bm{x}^{k}-\gamma_{k}^{-1}A^{\top}\bm{z}$. The detailed derivation of the dual problem (5.4) is relegated to Appendix A.1. From the property of the Moreau envelope of $\lambda\gamma_{k}^{-1}\|\cdot\|_{1}$ (see, e.g., [6, Proposition 12.29]), we see that $\Psi_{k}^{\text{reg}}$ is strongly convex and continuously differentiable with the gradient
@@
It is clear that all elements in ^$\widehat{\partial}^{2}\Psi_{k}^{\text{reg}}(\bm{z})$ are positive definite. We are now ready to present the Ssn method for solving equation (5.5) in Algorithm 2 and refer readers to [24, Theorem 3.6] for its convergence results.
@@
We next show that our inexact stopping criteria (SC1) and (SC2) can be achieved through some appropriate manipulations based on the dual sequence generated by the Ssn method. Specifically, at the $k$-th iteration, we apply the Ssn method for solving equation (5.5), which generates a dual sequence $\{\bm{z}^{k,t}\}$. Let
@@ 5.1.2 Comparison results
We also include the general iterative shrinkage and thresholding (GIST) algorithm [16], the non-monotone accelerated proximal gradient method with line search (denoted by nmAPG) [^1] The implementations of nmAPG in our experiments are based on the original Matlab codes, which are available at https://zhouchenlin.github.io/. [23], and the proximal difference-of-convex algorithm with extrapolation (pDCAe) [^2] The Matlab codes of the pDCAe for solving (5.1) are available at https://www.polyu.edu.hk/ama/profile/pong/pDCAe_final_codes/. [50] in our comparisons. Note that pDCAe can be viewed as a special case of Bregman proximal DC algorithm with extrapolation (BPDCAe) in [45], where the corresponding kernel function is chosen as $\phi(\bm{x})=\frac{1}{2}\|\bm{x}\|^{2}$. For each of these algorithms, we follow the parameter settings recommended in their respective references.
@@
In the following experiments, we choose $\lambda\in\{0.01,0.1,1,10\}$ and consider $(m,n,s)=(100i,1000i,20i)$ for $i=2,5,10,15,20$. For each triple $(m,n,s)$, we follow [50, Section 5] to randomly generate a trial as follows. First, we generate a matrix $A\in\mathbb{R}^{m\times n}$ with i.i.d. standard Gaussian entries. We then choose a subset $\mathcal{S}\subset\{1,\cdots,n\}$ of size $s$ uniformly at random and generate an $s$-sparse vector $\bm{x}_{\text{orig}}\in\mathbb{R}^{n}$, which has i.i.d. standard Gaussian entries on $\mathcal{S}$ and zeros on the complement set $\mathcal{S}^{c}$. Finally, we generate the vector $\bm{b}\in\mathbb{R}^{m}$ by setting $\bm{b}=A\bm{x}_{\text{orig}}+0.01\cdot\widehat{\bm{n}}$, where $\widehat{\bm{n}}\in\mathbb{R}^{m}$ is a random vector with i.i.d. standard Gaussian entries.
@@
The average computational results for each triple $(m,n,s)$ from 20 instances are presented in Tables 1 and 2. From the results, one can observe that both iBPDCA-SC1 and iBPDCA-SC2 exhibit superior numerical performance compared to GIST, nmAPG, and pDCAe, especially when the regularization parameter $\lambda$ is small or the problem size is large. For example, for the cases where $\lambda\leq 1$ and $n\geq 10000$, both iBPDCA-SC1 and iBPDCA-SC2 are consistently about 8 times faster than GIST, 5 times faster than pDCAe, and 2 times faster than nmAPG, while achieving comparable or even better objective function values. Moreover, one can also see that iBPDCA-SC1 and iBPDCA-SC2 exhibit the similar performance. This is indeed expected since they essentially employ the same framework, but differ in their stopping criteria for solving the subproblems. With the kernel function chosen as $\phi(\bm{x})=\frac{1}{2}\|\bm{x}\|^{2}$ in this part of experiments, the cost of computing the associated Bregman distance becomes negligible, rendering the effort required to verify inequalities (5.6) and (5.7) almost identical. Therefore, the CPU time of iBPDCA-SC1 and iBPDCA-SC2 can be similar when they undergo a similar number of total Ssn iterations.
@@
We also test all the algorithms with two instances $(A,\bm{b})$ obtained from the data sets mpg7 and gisette in the UCI data repository [20]. Here, the mpg7 data set is an expanded version of its original version, with the last digit signifying that an order 7 polynomial is used to generate the basis functions; see [24, Section 4.1] for more details. For mpg7, the matrix $A$ is of size $392\times 3432$ with $\lambda_{\max}(A^{\top}A)\approx 1.28\times 10^{4}$. For gisette, the matrix $A$ is of size $1000\times 4971$ with $\lambda_{\max}(A^{\top}A)\approx 3.36\times 10^{6}$. For each data set, we choose $\lambda=\lambda_{c}\|A^{\top}\bm{b}\|_{\infty}$ with $\lambda_{c}\in\{10^{-3},10^{-4},10^{-5}\}$. Figure 1 shows the numerical results of GIST, nmAPG, pDCAe, iBPDCA-SC1 and iBPDCA-SC2, where we plot the objective value $F_{\text{reg}}(\bm{x}^{k})$ against the computational time. From the results, we see that GIST, nmAPG, and pDCAe generally exhibit a slower rate of reduction in the objective value, potentially attributed to the large Lipschitz constant associated with the real data. In contrast, our iBPDCA-SC1 and iBPDCA-SC2 continue to demonstrate superior performance on these two real data sets.
@@
$\quad&\hskip 79.6678pt{\text{pDCAe}}\\ &\hskip 91.04872pt\Downarrow\\ &\min\limits_{\bm{x}\in\mathbb{R}^{n}}~{}{\overbrace{{\lambda\|\bm{x}\|_{1}}}^{P_{1}(\bm{x})}\,-{\overbrace{{\lambda\|\bm{x}\|}}^{P_{2}(\bm{x})}}+\,{\overbrace{{\frac{1}{2}\|A\bm{x}-\bm{b}\|_{2}^{2}}}^{f(\bm{x})}}}\end{aligned}\quad~{}~{}vs~{}~{}\quad\begin{aligned} &\hskip 79.6678pt{\text{iBPDCA}}\\ &\hskip 91.04872pt\Downarrow\\ &\min\limits_{\bm{x}\in\mathbb{R}^{n}}~{}{\overbrace{{\lambda\|\bm{x}\|_{1}+\frac{1}{2}\|A\bm{x}-\bm{b}\|_{2}^{2}}}^{P_{1}(\bm{x})}\,-\overbrace{{\lambda\|\bm{x}\|}}^{P_{2}(\bm{x})}\,+\,\overbrace{{0}}^{f(\bm{x})}}.\end{aligned}$
These two decomposition schemes result in different algorithms with different numerical performances. Since pDCAe also approximates the smooth convex part $\frac{1}{2}\|A\bm{x}-\bm{b}\|^{2}$ by a quadratic majorant, its performance would be significantly influenced by the Lipschitz constant $L_{A}:=\lambda_{\max}(A^{\top}A)$. This sensitivity to the Lipschitz constant is indeed a well-recognized limitation of first-order methods and may lead to the inferior performance of GIST, nmAPG and pDCAe for solving large-scale problems, as observed in Tables 1&\&&2 and Figure 1. This is because, for a randomly generated matrix $A$, the largest eigenvalue of $A^{\top}A$ tends to increase with the size, and for the data sets mpg7 and gisette, the largest eigenvalue of $A^{\top}A$ is around $1.28\times 10^{4}$ and $3.36\times 10^{6}$, respectively. In contrast, our iBPDCA can sidestep the majorization of the least squares term and choose instead to solve the subproblem via a highly efficient second-order Ssn method. The results in Tables 1&\&&2 and Figure 1 illustrate the promising performance of this algorithmic framework. But we should emphasize that these achievements are made possible by allowing the inexact minimization of the subproblem. This also highlights the motivation for developing an efficient inexact algorithmic framework in this work.
@@
(200,2000,40) GIST 3.22e-1 30000 2.74 0.03 2.54e+0 16498 1.45 0.03
nmAPG 2.54e-1 4507 0.40 0.03 2.54e+0 1535 0.14 0.03
pDCAe 2.73e-1 30000 2.06 0.03 2.54e+0 8353 0.56 0.03
iBPDCA-SC1 2.54e-1 120 (405) 0.64 0.03 2.54e+0 30 (170) 0.29 0.03
iBPDCA-SC2 2.54e-1 120 (414) 0.64 0.03 2.54e+0 30 (173) 0.30 0.03
(500,5000,100) GIST 9.92e-1 30000 14.47 0.16 7.02e+0 29927 14.28 0.16
nmAPG 6.92e-1 7634 3.49 0.16 6.92e+0 2660 1.29 0.16
pDCAe 8.12e-1 30000 10.60 0.16 6.92e+0 24732 8.62 0.16
iBPDCA-SC1 6.92e-1 161 (599) 5.10 0.16 6.92e+0 35 (259) 2.66 0.16
iBPDCA-SC2 6.92e-1 161 (612) 5.17 0.16 6.92e+0 35 (263) 2.68 0.16
(1000,10000,200) GIST 2.39e+0 30000 212.54 1.92 1.59e+1 30000 211.05 1.91
nmAPG 1.48e+0 11233 69.61 1.92 1.48e+1 4356 34.83 1.91
pDCAe 1.88e+0 30000 137.02 1.92 1.50e+1 30000 137.30 1.91
iBPDCA-SC1 1.48e+0 242 (898) 31.63 1.92 1.48e+1 50 (379) 16.71 1.91
iBPDCA-SC2 1.48e+0 242 (914) 31.86 1.92 1.48e+1 50 (382) 16.52 1.91
(1500,15000,300) GIST 3.86e+0 30000 552.27 5.09 2.53e+1 30000 547.84 5.09
nmAPG 2.24e+0 13908 217.94 5.09 2.24e+1 5412 114.91 5.09
pDCAe 2.99e+0 30000 362.98 5.09 2.31e+1 30000 363.23 5.09
iBPDCA-SC1 2.24e+0 299 (1072) 86.72 5.09 2.24e+1 59 (439) 44.33 5.09
iBPDCA-SC2 2.24e+0 299 (1088) 87.13 5.09 2.24e+1 59 (444) 44.19 5.09
(2000,20000,400) GIST 5.31e+0 30000 1018.43 9.35 3.49e+1 30000 1005.39 9.34
nmAPG 3.00e+0 16571 459.36 9.35 3.00e+1 6394 242.11 9.34
pDCAe 4.12e+0 30000 653.66 9.35 3.13e+1 30000 651.64 9.34
iBPDCA-SC1 3.00e+0 367 (1295) 201.98 9.35 3.00e+1 73 (518) 98.71 9.34
iBPDCA-SC2 3.00e+0 367 (1312) 202.73 9.35 3.00e+1 73 (524) 98.39 9.34
@@
(200,2000,40) GIST 2.53e+1 704 0.07 0.03 2.44e+2 94 0.01 0.03
nmAPG 2.53e+1 301 0.03 0.03 2.44e+2 161 0.02 0.03
pDCAe 2.53e+1 515 0.04 0.03 2.44e+2 153 0.01 0.03
iBPDCA-SC1 2.53e+1 9 (50) 0.10 0.03 2.44e+2 8 (32) 0.08 0.03
iBPDCA-SC2 2.53e+1 9 (50) 0.10 0.03 2.44e+2 8 (32) 0.08 0.03
(500,5000,100) GIST 6.90e+1 4434 2.22 0.16 6.80e+2 240 0.15 0.16
nmAPG 6.90e+1 683 0.41 0.16 6.80e+2 198 0.12 0.16
pDCAe 6.90e+1 2633 0.94 0.16 6.80e+2 237 0.10 0.16
iBPDCA-SC1 6.90e+1 12 (116) 1.29 0.16 6.80e+2 7 (45) 0.57 0.16
iBPDCA-SC2 6.90e+1 12 (117) 1.30 0.16 6.80e+2 7 (45) 0.58 0.16
(1000,10000,200) GIST 1.48e+2 14834 105.07 1.91 1.47e+3 585 4.33 1.90
nmAPG 1.48e+2 1421 14.14 1.91 1.47e+3 268 1.86 1.90
pDCAe 1.48e+2 8547 39.06 1.91 1.47e+3 401 1.87 1.90
iBPDCA-SC1 1.48e+2 15 (185) 9.74 1.91 1.47e+3 7 (69) 4.11 1.90
iBPDCA-SC2 1.48e+2 15 (184) 9.64 1.91 1.47e+3 7 (68) 4.10 1.90
(1500,15000,300) GIST 2.24e+2 23697 434.23 5.09 2.23e+3 1213 22.69 5.09
nmAPG 2.24e+2 2014 56.34 5.09 2.23e+3 409 9.10 5.09
pDCAe 2.24e+2 13670 165.42 5.09 2.23e+3 777 9.52 5.09
iBPDCA-SC1 2.24e+2 17 (225) 27.07 5.09 2.23e+3 8 (93) 13.07 5.09
iBPDCA-SC2 2.24e+2 17 (226) 27.00 5.09 2.23e+3 8 (93) 12.98 5.09
(2000,20000,400) GIST 2.99e+2 28289 952.63 9.32 2.98e+3 2661 91.11 9.29
nmAPG 2.99e+2 2506 128.11 9.32 2.98e+3 623 29.32 9.29
pDCAe 2.99e+2 20631 448.91 9.32 2.98e+3 1685 37.01 9.29
iBPDCA-SC1 2.99e+2 20 (264) 59.98 9.32 2.98e+3 9 (129) 33.45 9.29
iBPDCA-SC2 2.99e+2 20 (265) 59.49 9.32 2.98e+3 9 (129) 33.23 9.29
@@
Figure 1: Numerical results of GIST, nmAPG, pDCAe, iBPDCA-SC1 and iBPDCA-SC2 for solving the $\ell_{1-2}$ regularized least squares problem on mpg7 and gisette from the UCI data repository.